the length of a room exceeds its breath by 3dm. If the length of a room had been increased by 3dm and breath is diminished by 2dm, the area would have not been altered. Find the dimensions.
Answer by ikleyn(52787)

.
the length of a room exceeds its breath by 3dm. If the length of a room had been increased by 3dm and breath is diminished by 2dm,
the area would have not been altered. Find the dimensions.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

x - y = 3,                (1)
(x+3)*(y-2) - xy = 0.     (2)

Simplify (by opening parentheses in (2) )

x - y = 3,
xy +3y - 2x -6 - xy = 0,

or

  x -  y = 3,             (1')
-2x + 3y = 6.             (2')

Express x = 3+y from (1') and substitute it into (2'). You will get

-2(3+y) + 3y = 6.

Single equation for one unknown.

I am sure that you can complete the solution from this point.